One More Time Software has 7.2 percent coupon bonds on the market with 17 years to maturity. The bonds make semiannual payments and currently sell for 84.5 percent of par.

Required:
(a) What is the current yield on the bonds?

(Click to select)

8.95%

0.09%

8.52%

0.19%

9.45%

(b) The YTM?

(Click to select)

8.87%

9.18%

7.70%

4.50%

9.00%

(c) The effective annual yield?

(Click to select

)9.20%

8.74%

0.09%

9.66%

8.97%
Answer #1

a) Current yield = Coupon payment / Current price = ( 1000 * 7.2% ) / 845 = 8.52%

b) YTM = 9%

Calculated as below in excel or financial calculator:

NPER 34 [17*2]
FV 1000
PMT 36.00 [100*7.2%*1/2]
PV 845.0
Current yield 4.50% [Rate ( nper, pmt,-pv,fv)]
YTM 9.00% [4.5%*2]

c) Effective annual yield = ( 1 + YTM / 2) 2 -1

= ( 1+ 9% /2 ) 2 -1

= 9.20%